OpenStudy (anonymous):

What is the slope of the line that passes through (1, 4) and (1, –3)?

OpenStudy (anonymous):

appy two point form formula............

OpenStudy (anonymous):

question makes no sense since x does not increase from 1 to 1 slope is undefined

OpenStudy (anonymous):

Since x is same for both points, hence it is a vertical line. Hence, the slope of this line is infinity

OpenStudy (anonymous):

argh

OpenStudy (anonymous):

actaully it is not infinity

OpenStudy (anonymous):

parallel to y axis.

razor99 (razor99):

fool a little help plzzzzzzzzzzzzz.

OpenStudy (anonymous):

it is just not a number. don't let anyone tell you that \[\frac{1}{0}=\infty\] it is not true, not even as a limit statement (which it is not)

OpenStudy (anonymous):

No you are correct, I admit

OpenStudy (anonymous):

satelite73 i love you dude #nohomo

OpenStudy (anonymous):

we should say undefined, but the question makes sense :)

OpenStudy (anonymous):

i see the question often enough to know it is frequently asked, but it is akin to asking what the gas milage is on my bicycle, so in fact it doesn't make much sense

OpenStudy (anonymous):

what is \( \infty \)?

OpenStudy (anonymous):

Sorry, not in much of mood for argument, but these two are not same, and finally the question seemingly now making some sense